If P(A) = 0.7, P(B) = 0.6, and P(A or B) = 0.8, find P(A and B).
P(A and B) =

Answer by ikleyn(52747)

.

Use one of the basic identities of the Probability theory

    P(A or B) = P(A) + P(B) - P(A and B).


Substitute the given date into the identity. You will get

    0.8 = 0.7 + 0.6 - P(A and B).


It implies

    P(A and B) = 0.7 + 0.6 - 0.8 = 0.5.    ANSWER

-------------

Solved, answered and explained in all details.